OpenStudy (anonymous):

Plzz help! Choose the system of equations which matches the following graph.

OpenStudy (kidrah69):

following..........?

OpenStudy (anonymous):

here's the graph for the Q ^ @kidrah69 @They_Call_Me_Narii

OpenStudy (anonymous):

A. x + 2y = −10 3x − y = 12 B. x − 2y = −10 3x − y = 12 C. x + 2y = −10 3x + y = 12 D. x − 2y = −10 3x + y = 12

OpenStudy (sparrow2):

it's a just use graphs

OpenStudy (anonymous):

@CallMeKiki

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

any guesses? @They_Call_Me_Narii

OpenStudy (anonymous):

No. . I Dont Wanna Give U The Wrong Answer

OpenStudy (anonymous):

well if you had a guess what would it be? @They_Call_Me_Narii

OpenStudy (anonymous):

@alyssa_xo can you help?

OpenStudy (alyssa_xo):

the solution is the point where the graphs are equal the solution (graphically) is the point where the lines intersect find the point where (2,-6) satisfies both equations

OpenStudy (alyssa_xo):

also worth mentioning the line with a positive slope has a slope of 3 we can cancel c and d because when we get the line into y=mx+b form, our slope is -3 \(3x + y = 12 \\y=\color{red}{-3}x+12\)

OpenStudy (anonymous):

ok I think it's either A or D am I right?

OpenStudy (alyssa_xo):

mhm

OpenStudy (anonymous):

So would it be A ?

OpenStudy (alyssa_xo):

yeah. check your work by plugging in the point of intersection (i.e. the point on both lines)

OpenStudy (anonymous):

ok thx for your help :) ^

OpenStudy (alyssa_xo):

np np